Suppose a chemist combines a 25% acid solution and a 50% acid solution to make 40 l of 45% acid solution. How many liters of each solution did she use?
Answer by stanbon(75887)

Suppose a chemist combines a 25% acid solution and a 50% acid solution to make 40 l of 45% acid solution. How many liters of each solution did she use?
------
acid + acid = acid
----------------------
0.25x + 0.50(40-x) = 0.45*40
------------------------
25x + 50*40 - 50x = 45*40
-25x = -5*40
x = (1/5)40
x = 8 liters (amt of 25% solution needed)
40-x = 32 liters (amt. of 50% solution needed)
=====================
Cheers,
Stan H.
=====================